THE ASTROCHYMIST: AN INTERNET RESOURCE FOR ASTROCHEMISTS AND INTERESTED BY-STANDERS

\underline{DAVID E. WOON}, Molecular Research Institute, 2495 Old Middlefield Way, Mountain View, CA 94043.

The objective of The Astrochymist website is to provide content for the world-wide community of professionals active in the field of astrochemistry as well as non-professionals who are interested in the subject. The initial content includes a frequently-updated list of publications of interest to astrochemists, a top-level treatment of Frequently Asked Questions (FAQ), listings of journals, conferences, and people active in the field of astrochemistry, plus the beginnings of a comprehensive bibliography of astromolecules grouped by the environment in which they are found (interstellar & circumstellar media, comets, planetary & satellite atmospheres, and brown dwarfs & stellar atmospheres). One molecule is designated Astromolecule of the Month. Suggestions for new types of content will be very welcome.